# Efficient Action Counting with Dynamic Queries

### v2

@@ 3.3 Dynamic Action Query
As discussed in Sec. 1, the TRC problem requires recognizing open-set action instances depending on the video content, where the action category is not predefined. Therefore, we propose the Dynamic Action Query strategy, which adaptively updates the content query $\mathbf{Q}^{\text{cnt}}$ using distilled content features from the encoder. Simple yet effective, the DAQ strategy directly initializes the input content queries $\mathbf{Q}^{\text{cnt}}$ of the decoder using the $Q$ content queries filtered by the query selection module, as shown in Fig. 2. This mechanism provides priors for the decoder that allows it to attend to the “action of interest” conditioned on the input video contents in a dynamic and contextually aware manner. In this way, we avoid the need for the manual definition of the action categories, thereby improving the models’ generalization capability.
Furthermore, we extensively explore different methods for initializing the queries in the decoder $\mathcal{D}(\cdot)$ in the supplementary material, confirming the DAQ strategy as the most effective.
@@ 3.4 Inter-query Contrastive Learning
Since the input video may contain multiple action categories and other distractors such as the background motion, another unique challenge to the TRC task is to recognize action instances with identical content. This requires that the actions of interest we classify exhibit similarity in their motion patterns, while other action queries should have dissimilar representations. To tackle this challenge, we propose Inter-query Contrastive Learning to distinguish the action queries. Intuitively, we hope to partition the content queries into $K=2$ main categories through contrastive learning. One category comprises the “actions of interest”, with similar representations, while the other consists of null instances. The latter may encompass interfering actions or background elements, with distinct representations from those of the actions of interest.
Concretely, we designate the $Q$ decoder features $\mathbf{H}^{\text{dec}}$ features classified into two sets according to the predicted class labels $\mathbf{C}^{\text{cls}}$, as shown in Fig. 2. The features that are classified as actions of interest form the positive set $S^{+}$, while the remaining features form the negative set $S^{-}$. Then we apply contrastive learning using InfoNCE loss [12] over the feature space:
$\begin{split}\mathcal{L}_{+}&=\sum\limits_{s\in S^{+},s\neq q}{\rm exp}(\mathbf{H}^{\text{dec}}_{q}\cdot\mathbf{H}^{\text{dec}}_{s})/\tau,\quad\mathcal{L}_{-}=\sum\limits_{s\in S^{-}}{\rm exp}(\mathbf{H}^{\text{dec}}_{q}\cdot\mathbf{H}^{\text{dec}}_{s})/\tau,\\ \mathcal{L}_{\text{ctrs}}&=-\sum\limits_{q\in S^{+}}{\rm log}\left(\frac{\mathcal{L}_{+}}{\mathcal{L}_{+}+\mathcal{L}_{-}}\right),\end{split}$ (3)
@@
3.5 Training Loss
The overall loss function is defined as:
$\mathcal{L}=\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}+\lambda_{\text{pos}}\mathcal{L}_{\text{pos}}+\lambda_{\text{ctrs}}\mathcal{L}_{\text{ctrs}},$ (4)
where $\lambda_{\text{cls}}$, $\lambda_{\text{pos}}$, $\lambda_{\text{ctrs}}$ are the coefficients of each loss term, respectively.
@@ 4.1 Datasets and Metrics
RepCountA dataset [13] is currently the largest and most challenging benchmark for the video TRC task  [^1] The RepCountB [13] test subset is proprietary and not publicly available.. It is primarily compiled from fitness videos on YouTube, including a wide range of fitness activities conducted in diverse settings, including homes, gyms, and outdoor environments. This dataset stands out due to its extensive video lengths, significant variations in the average motion cycle, and a higher number of repetitive cycles compared to prior datasets [9, 32, 17, 46]. We use the start and end positions of each action instance $\hat{\mathbf{P}}_{q}$ provided by the annotations, and disregard the class action labels. We train our model on the RepCountA train set and select the best model on the validation set. We report the evaluation results on the test set.
Metrics. We compute two commonly used metrics, OBO and MAE  [13, 9, 46], to evaluate the model performance. OBO (Off-By-One count error) measures the percentage that the predicted count is within the GT count ±$\pm 1$ range. MAE (Mean Absolute Error) measures the normalized absolute difference between the predicted and GT counts. Formally,
${\rm OBO}=\frac{1}{M}\sum_{i=1}^{M}\left|N_{i}-\hat{N}_{i}\leq 1\right|,\quad{\rm MAE}=\frac{1}{M}\sum_{i=1}^{M}\frac{\left|N_{i}-\hat{N}_{i}\right|}{N_{i}},$ (5)
where ${N}_{i}$ and $\hat{N}_{i}$ are the predicted and GT counts for the $i^{th}$ test video, respectively, and $M$ is the total number of test videos.
4.2 Implementation Details
We implement our approach with different backbone video feature extractors respectively, including TSN [40], I3D [4], and ViT-B [8]. For the Transformer architecture, we employ a 2-layer encoder and a 4-layer decoder, both with 8-head attention mechanisms. The feature dimension is set to $C=512$. We set $Q=40$ queries in our method empirically. For more results with different feature channels $C$ and query numbers $Q$, please refer to the supplementary materials. The length of the video input is set to $T=512$ frames without down-sampling. We utilize the AdamW optimizer [26] with a learning rate of $0.002$, a batch size of $64$, and train the model for $80$ epochs. We set $\lambda_{cls}=1.0$, $\lambda_{pos}=5.0$, $\lambda_{gIoU}=0.4$, $\lambda_{ctrs}=1.0$, confidence threshold $\alpha=0.2$. We provide further implementation details in the supplementary.
4.3 Comparison to State-of-the-arts
Table 1: Comparison to the state-of-the-arts on RepCountA [13] dataset. We compare with SOTA action recognition/segmentation methods (top block), TRC methods (second block), and action detection methods (third block with normal-$\dagger$). We further report MAE and OBO metrics for short-, medium-, and long-period test actions.
@@ 4.3 Comparison to State-of-the-arts
RepNet [9] ResNet-50 0.5865 0.2450 0.7793 0.0930 0.5893 0.1591 0.4549 0.4062
Zhang et al. [46] 3D-ResNext101 0.8786 0.1554 - - - - - -
TransRAC [13] Swin-T 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
Li et al. [18] Swin-T 0.3841 0.3860 - - - - - -
@@
Ours ViT-B 0.2622 0.5430 0.2257 0.2558 0.2002 0.5909 0.3294 0.7031
As shown in Tab. 1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths. Specifically, the similarity-matrix-based methods [9, 13, 18] suffer from quadratic computation complexity. To manage this, they employ a sparse sampling strategy, ensuring reasonable content coverage within a limited temporal context window. However, this approach results in inferior performance for short, rapid action instances, as it tends to overlook the cycles. Conversely, the action detection approaches [23, 44, 33] are primarily developed for detecting action instances specific to particular classes. To adapt them to the TRC task, we modify their output layer to incorporate class-agnostic supervision accordingly. However, relying solely on class-agnostic supervision does not provide them with the ability to dynamically identify repetitive cycles based on the input videos. As a result, their performance on long, slow action instances is generally inferior compared to dedicated TRC approaches. In contrast, our approach effectively balances the detection of actions at various speeds.
@@
RepNet [9] ResNet-50 0.5336 0.2984 0.6219 0.1739 0.4825 0.3600 0.4996 0.5000
TransRAC [13] Swin-T 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
Li et al. [18] 3D-ResNext101 0.5227 0.3500 - - - - - -
@@
4.4 Qualitative Results
We visualize the predictions of our approach and baseline method TransRAC [13] on the RepCountA [13] dataset to better demonstrate our performance improvement in Fig. 3. TransRAC [13] represents action cycles using density maps, and the final count value is obtained by summing the values in the density map. However, this approach suffers from a lack of interpretability and finally results in miscounting. In contrast, our approach not only produces an accurate final count but also precisely localizes the action start and end positions (colored blocks) in most cases, aligning with the GT labels. It can be seen that our method exhibits robustness to changes in viewpoint, background noise, and sudden interruptions. For instance, in the second case of Fig. 3, we accurately and robustly estimate the time positions even with viewpoint changes as well as a sudden interruption in the middle of the timeline.
@@ 4.4 Qualitative Results
We further illustrate the generalization performance of the proposed method on the unseen UCFRep [46] test set in Fig. 5. We directly apply our trained model and do not use UCFRep training data. Our model still accurately recognizes the action instances and gets the correct count in the challenging cases, indicating robust generalization ability. Specifically, the top case exhibits extreme viewpoint and lighting conditions, while the bottom case contains the action of soccer juggling which is not seen in the training set. We attribute the performance advantage to the proposed DAQ and ICL designs, which empower the model to adaptively adjust the action queries based on the input video features and effectively localize similar (repetitive) action instances, distinguishing them from the background noise actions. Please refer to our supplementary video for more qualitative results which are robust to varying cycle lengths, sudden interruptions, and multiple distracting action classes.
@@
4.5 Ablation Study
Effect of DAQ and ICL. We implement two ablated models to study the efficacy of the proposed DAQ and ICL designs. Tab. 4 presents the results on RepCountA [13] using the TSN backbone. In ablation (a), we substitute the proposed DAQ module with a static content query, where the content queries $\mathbf{Q}^{\text{cnt}}$ are also learnable variables, instead of using the content queries filtered out by the encoder and query selection modules. In ablation (b), we eliminate the ICL design among the action queries. The results demonstrate that both DAQ and ICL contribute to performance improvement, particularly in terms of counting medium and long actions.
@@ 4.5 Ablation Study
Confidence threshold $\alpha$. We evaluate the impact of different confidence thresholds for the “action of interest” on the final counting results in Fig. 6 (green curve). The results indicate that setting the threshold within the range of $0.2$ to $0.4$ yields similar performance for our method. Our method consistently outperforms TransRAC [13] (blue curve) by a considerable margin.
Figure 6: Results of different confidence thresholds of our method and TransRAC [13]. We depict the MAE (left) and OBO (right) curves of our method (green curve) and the TransRAC (blue curve) approach with regard to different confidence thresholds. The metrics of TransRAC are obtained by binarizing the density map of TransRAC output and then summing to obtain the final count.
@@ 5 Conclusion
In conclusion, our study introduces an innovative perspective for the TRC task, which reduces the computational complexity and maintains robustness across varying action periods and video lengths. To tackle the challenge of open-set action categories, we propose DAQ, which improves generalization across different actions. To recognize actions with identical content, we propose ICL, which facilitates learning repetitive action representations and distinguishing them from distractions. Integrating DAQ and ICL, our method can identify contextually similar action instances that are adaptively conditioned on the video content. Experimental results on challenging benchmarks demonstrate the superiority of our approach compared to SOTAs, in terms of both accuracy and efficiency. Furthermore, our approach adeptly balances the demands of diverse action speeds and video durations, establishing a solid foundation for practical implementations in real-world scenarios.

### v3

@@ 3.4 Dynamic Action Query
As discussed in Sec. 1, the TRC problem requires recognizing open-set action instances depending on the video content, where the action category is not predefined. Therefore, we propose the Dynamic Action Query strategy, which adaptively updates the action query $\mathbf{Q}^{\text{act}}$ using the selected encoder action query features $\mathbf{E}^{\text{act}}$, i.e. $\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$, as shown in Fig. 2. The DAQ strategy, simple yet effective, provides priors for the decoder to dynamically and contextually attend to “repetitive actions” based on the input video content, eliminating the need for manually defining action categories and thereby enhancing the model’s generalization capability. This mechanism ensures that the cross-attention module in the decoder pools features from the entire video while focusing on relevant local repetitive features. Such dynamic adjustment allows the model to concentrate on pertinent “repetitive actions”, offering a deeper understanding of action queries and improving training convergence.
In summary, the DAQ approach not only circumvents the need for manual action category definitions but also leverages dynamic, contextually aware priors to enhance both the focus and generalization capabilities of the model. Our supplementary material further explores various methods for initializing the queries in the decoder $\mathcal{D}(\cdot)$, confirming that the DAQ strategy is the most effective.
@@ 3.5 Inter-query Contrastive Learning
Since the input video may contain other distractors such as the background motion  [^2] We assume that each video contains only a primary repetitive action type, which is also the case for the public dataset., another unique challenge to the TRC task is to recognize action instances with identical content. This requires that the actions of interest we classify exhibit similarity in their motion patterns, while other action queries should have dissimilar representations. To tackle this challenge, we propose Inter-query Contrastive Learning to distinguish the action queries. Intuitively, based on the classification predictions $\mathbf{A}$, we partition the action queries decoded by the decoder $\mathbf{D}^{\text{act}}$ into two categories and employ inter-query contrastive learning on them, as shown in Fig. 2 (right). The features that are classified as “repetitive actions” form the positive set $S^{+}$, while the other features form the negative set $S^{-}$. Then we apply contrastive learning using InfoNCE loss [12] $\mathcal{L}_{\text{ctrs}}$ over the representation space:
$\begin{split}\mathcal{L}_{\text{ctrs}}&=-\sum\limits_{i\in S^{+}}{\rm log}\left(\frac{\mathcal{L}_{+}}{\mathcal{L}_{+}+\mathcal{L}_{-}}\right),\\ \mathcal{L}_{+}&=\sum\limits_{s\in S^{+},s\neq i}{\rm exp}(\mathbf{D}^{\text{act}}_{i}\cdot\mathbf{D}^{\text{act}}_{s})/\tau,\\ \mathcal{L}_{-}&=\sum\limits_{s\in S^{-}}{\rm exp}(\mathbf{D}^{\text{act}}_{i}\cdot\mathbf{D}^{\text{act}}_{s})/\tau,\end{split}$ (5)
@@
3.6 Training
We train our model in an end-to-end manner using the overall loss function:
$\mathcal{L}=\lambda_{\text{Hungarian}}\mathcal{L}_{\text{Hungarian}}+\lambda_{\text{ctrs}}\mathcal{L}_{\text{ctrs}},$ (6)
where $\lambda_{\text{Hungarian}},\lambda_{\text{ctrs}}\in\mathbb{R}$ are the coefficients. Following DETR [3], we also found it helpful to use auxiliary losses in the decoder during training. Specifically, we add $\mathcal{L}_{\text{Hungarian}}$ loss on the predictions from the prediction heads right after the encoder i.e. $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{P}}$, and add prediction heads and $\mathcal{L}_{\text{Hungarian}}$ loss after each decoder layer. All prediction heads share their parameters. This ensures the model focuses on the correct features at each stage consistently, thereby accelerating convergence.
@@ 4.1 Datasets and Metrics
RepCountA dataset [14] is currently the largest and most challenging benchmark for the video TRC task  [^3] The RepCountB [14] test subset is proprietary and not publicly available.. It is primarily compiled from fitness videos on YouTube, including a wide range of fitness activities conducted in diverse settings, including homes, gyms, and outdoor environments. This dataset stands out due to its extensive video lengths, significant variations in the average motion cycle, and a higher number of repetitive cycles compared to prior datasets [9, 33, 18, 47]. We use the start and end positions of each action instance to compute the GT label $\hat{\mathbf{P}}$ provided by the annotations. We train our model on the RepCountA train set and select the best model on the validation set. We report the evaluation results on the test set.
Table 1: Comparison to the state-of-the-arts on RepCountA [14] dataset. We compare with SOTA action recognition/segmentation methods (top block), TRC methods (second block), and action detection methods (third block with $\dagger$). We further report MAE and OBO metrics for short-, medium-, and long-period test actions.
@@
RepNet [9] ResNet 0.5865 0.2450 0.7793 0.0930 0.5893 0.1591 0.4549 0.4062
Zhang et al. [47] 3D-ResNext 0.8786 0.1554 - - - - - -
TransRAC [14] ViT 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
Li et al. [19] ViT 0.3841 0.3860 - - - - - -
@@
Ours ViT 0.2622 0.5430 0.2257 0.2558 0.2002 0.5909 0.3294 0.7031
Metrics. Following previous works [14, 9, 47, 19], we compute two commonly used metrics, OBO and MAE, to evaluate the model performance. OBO (Off-By-One count error) measures the percentage that the predicted count is within the GT count ±$\pm 1$ range. MAE (Mean Absolute Error) measures the normalized absolute difference between the predicted and GT counts. Formally,
${\rm OBO}$ $=\frac{1}{M}\sum_{m=1}^{M}\left|N_{m}-\hat{N}_{m}\leq 1\right|,$ (7)
${\rm MAE}$ $=\frac{1}{M}\sum_{m=1}^{M}\frac{\left|N_{m}-\hat{N}_{m}\right|}{N_{m}},$ (8)
where $M$ is the total number of test videos, ${N}_{m}$ and $\hat{N}_{m}$ are the predicted and GT counts for the $m^{th}$ test video, respectively.
4.2 Implementation Details
We implement our approach with different backbone video feature extractors respectively, including TSN [41], I3D [4], and ViT [8]. For encoder-decoder Transformer, we set ${L}^{\text{enc}}=2$ and ${L}^{\text{dec}}=4$, both with 8-head attention mechanisms. The feature dimension is set to $C=512$. We set $Q=40$ queries in our method empirically. For more results with different feature channels $C$ and query numbers $Q$, please refer to the supplementary materials. The length of the video input is set to $T=512$ frames without down-sampling. We utilize the AdamW optimizer [27] with a learning rate of $0.002$, a batch size of $64$, and train the model for $80$ epochs. We set $\lambda_{\text{Hungarian}}=1.0$, $\lambda_{\text{L1}}=5.0$, $\lambda_{\text{gIoU}}=0.4$, $\lambda_{\text{ctrs}}=1.0$, confidence threshold $\alpha=0.2$. We provide further implementation details in the supplementary.
4.3 Comparison to State-of-the-arts
@@
RepNet [9] ResNet 0.5336 0.2984 0.6219 0.1739 0.4825 0.3600 0.4996 0.5000
TransRAC [14] ViT 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
Li et al. [19] 3D-ResNext 0.5227 0.3500 - - - - - -
@@
As shown in Tab. 1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths. Specifically, the similarity-matrix-based methods [9, 14, 19] suffer from quadratic computation complexity. To manage this, they employ a sparse sampling strategy, ensuring reasonable content coverage within a limited temporal context window. However, this approach results in inferior performance for short, rapid action instances, as it tends to overlook the cycles. Conversely, the action detection approaches [24, 45, 34] are primarily developed for detecting action instances specific to particular classes. To adapt them to the TRC task, we modify their output layer to incorporate class-agnostic supervision accordingly. However, relying solely on class-agnostic supervision does not support them to dynamically identify repetitive cycles based on the input videos. As a result, their performance on long, slow action instances is generally inferior compared to dedicated TRC approaches. In contrast, our approach effectively balances the detection of actions at various speeds.
@@
4.4 Qualitative Results
We visualize the predictions of our approach and baseline method TransRAC [14] on the RepCountA [14] dataset in Fig. 3. TransRAC [14] represents action cycles using density maps, and the final count value is obtained by summing the values in the density map. However, this approach suffers from a lack of interpretability and finally results in miscounting. In contrast, our approach not only produces an accurate final count but also correctly localizes the action start and end positions (colored blocks) in most cases. In addition, our method exhibits robustness to changes in viewpoint, background noise, and sudden interruptions, e.g., in the second case of Fig. 3, we accurately estimate the time positions even with viewpoint changes as well as a sudden interruption in the middle of the timeline. Please refer to our supplementary video for more qualitative results. We present the video result of the first case in Fig. 3, which shows that while the subject performs pull-ups, there are moments of talking and arm waving. Our method selectively counts only the pull-up actions, as they are the primary focus and consistent with the main action instances of the video, excluding dissimilar actions like “waving”, validating the robustness and generalization ability of our method.
We further illustrate the generalization performance of the proposed method on the unseen UCFRep [47] test set in Fig. 5. We directly apply our trained model and do not use UCFRep training data. Our model still accurately recognizes the action instances and gets the correct count in the challenging cases, indicating robust generalization ability. Specifically, the top case exhibits extreme viewpoint and lighting conditions, while the bottom case contains the action of soccer juggling which is not seen in the training set. We attribute the performance advantage to the proposed DAQ and ICL designs, which empower the model to adaptively adjust the action queries based on the input video features and effectively localize similar (repetitive) action instances, distinguishing them from the background noise actions.
4.5 Ablation Study
Effect of DAQ and ICL. We implement two ablated models to study the efficacy of the proposed DAQ and ICL designs. Tab. 4 presents the results on RepCountA [14] using the TSN backbone. In ablation (a), we substitute the proposed DAQ module with a static action query, where the action queries $\mathbf{Q}^{\text{act}}$ are also learnable variables, instead of using the action queries filtered out by the encoder and query selection modules. In ablation (b), we eliminate the ICL design among the action queries. The results demonstrate that both DAQ and ICL contribute to performance improvement, particularly in terms of counting medium and long actions.
Confidence threshold $\alpha$. We report the impact of different confidence thresholds for classifying “repetitive actions” on the final counts in Fig. 6. We illustrate the MAE (left) and OBO (right) curves for our method (green curve) and the TransRAC [14] approach (blue curve) across various confidence thresholds $\alpha$. The metrics for TransRAC are derived by binarizing its output density map and summing the results to obtain the final count. The results indicate that setting the threshold within the range of $0.2$ to $0.4$ yields similar performance for our method. Our method consistently outperforms TransRAC [14] (blue curve) by a considerable margin.
Figure 6: Results of different confidence thresholds of our method and TransRAC [14]. We depict the MAE (left) and OBO (right) curves of our method (green curve) and the TransRAC (blue curve) approach concerning different confidence thresholds. The metrics of TransRAC are obtained by binarizing the density map of TransRAC output and then summing to obtain the final count.
@@ 5 Conclusion
In conclusion, we provide an innovative perspective for tackling TRC tasks, which reduces computational complexity and maintains robustness across varying action lengths. To address open-set action categories, we propose DAQ for improved generalization, and ICL for recognizing repetitive actions and distinguishing them from distractions. Integrating DAQ and ICL, our method adaptively identifies contextually similar action instances. Experimental results on challenging benchmarks demonstrate our approach’s superiority over SOTAs in both accuracy and efficiency, while adeptly balancing diverse action speeds and video durations, establishing a solid foundation for practical implementations in real-world scenarios.
[…]
Acknowledgements This work was supported by National Science and Technology Major Project (2022ZD0114904).
[…]
Data Availability Statements. The RepCountA [14] and UCFRep [47] datasets that support the findings of this study are publicly available on GitHub at https://github.com/SvipRepetitionCounting/TransRAC and https://github.com/Xiaodomgdomg/Deep-Temporal-Repetition-Counting, respectively.